# Mixed-Integer Optimal Control via Reinforcement Learning: A Case Study on Hybrid Vehicle Energy Management

### v2

@@ Abstract
Many optimal control problems require the simultaneous output of continuous and discrete control variables. Such problems are usually formulated as mixed-integer optimal control (MIOC) problems, which are challenging to solve due to the complexity of the solution space. Numerical methods such as branch-and-bound are computationally expensive and unsuitable for real-time control. This brief proposes a novel continuous-discrete reinforcement learning (CDRL) algorithm, twin delayed deep deterministic actor-Q (TD3AQ), for MIOC problems. TD3AQ combines the advantages of both actor-critic and Q-learning methods, and can handle the continuous and discrete action spaces simultaneously. The proposed algorithm is evaluated on a plug-in hybrid electric vehicle (PHEV) energy management problem, where real-time control of the continuous variable, engine torque, and discrete variables, gear shift and clutch engagement/disengagement, is essential to maximize fuel economy while satisfying driving constraints. Simulation results on different drive cycles show that TD3AQ achieves near-optimal control compared to dynamic programming (DP) and outperforms baseline reinforcement learning algorithms.
Index Terms:
@@
Mixed-integer optimal control, reinforcement learning, hybrid action space, hybrid electric vehicle energy management.
I Introduction
MIXED-INTEGER optimal control (MIOC) problems commonly arise in various engineering applications, such as assignment problems [1], control of piecewise-affine (PWA) systems [2] and hybrid systems [3]. These problems are particularly challenging as they require optimization of a system that involves both continuous and discrete control variables.
One common approach for solving MIOC problems is to formulate them as mixed-integer programming (MIP) models and apply the branch-and-bound method [4]. This method involves dividing the solution space into smaller branches and using a bounding function to determine if each branch can be eliminated. Although it can provide globally optimal solutions, it is computationally expensive and time-consuming. Another approach is to use dynamic programming (DP), which breaks the problem into smaller sub-problems and solves them recursively [5]. However, DP requires discretization of continuous variables and knowledge of the entire state space in advance, making it less suitable for real-world optimal control. Heuristic algorithms offer feasible and fast solutions by leveraging problem-specific knowledge to guide the search [6, 7], but they may not provide optimality guarantees and require customization for each problem structure [7].
@@ I Introduction
In recent years, a substantial amount of RL research has focused on a special type of MIOC problems called parameterized action Markov decision process (PAMDP) [12]. The key of PAMDP is to formulate the problem as a hierarchical structure where each discrete action is associated with a set of continuous parameters. One approach is to alternate between optimizing the discrete action selection with Q-learning and the parameter selection with a policy search method [12]. Another approach called parameterized deep deterministic policy gradient (PA-DDPG) is to directly output the continuous parameters and weights for discrete actions from DDPG [13] and select the discrete action with the highest weight [14]. Parameterized deep Q-networks (P-DQN) combine the actor-critic architecture and DQN, where the actor network first outputs parameters for all discrete actions, and then the Q-network selects the action with the highest Q-value [15]. Other approaches utilize neural ordinary differential equations as state-value networks [16] or conditional variational auto-encoders [17] to handle the parameterized action space. Alternatively, some methods handle the mixed-integer action space in a unified manner. These approaches add extra layers [18] or parallel networks [19] to generate continuous and discrete actions, and then optimize the policy using techniques such as dueling DQN or proximal policy optimization.
In the domain of mixed-integer control for hybrid electric vehicle (HEV) energy management, previous studies have employed various approaches, including utilizing frameworks similar to PA-DDPG [20] or P-DQN [21]. Tang et al. [22] combined DQN and DDPG, obtaining the gear ratio and engine throttle independently. However, no studies used RL to concurrently tackle the optimization of gear shifting and engine on/off control, which represents an unexplored research area in this field.
@@
Inspired by the aforementioned research, this brief proposes a novel continuous-discrete reinforcement learning (CDRL) algorithm called twin delayed deep deterministic actor-Q (TD3AQ) for MIOC problems. The algorithm is applied to address an energy management problem for a plug-in hybrid electric vehicle (PHEV), aiming to demonstrate its effectiveness. The main contributions of this study are summarized as follows:
  1. 1)
    A state-of-the-art CDRL algorithm, TD3AQ, is proposed to address the mixed-integer action space. TD3AQ combines the advantages of both actor-critic and Q-learning methods, providing a unified solution for MIOC problems. Comparative evaluations demonstrate that TD3AQ achieves near-optimal solutions compared to dynamic programming (DP) and outperforms previous baselines [14, 23, 11]. Moreover, the sub-millisecond computation time of TD3AQ renders it suitable for real-time control applications.
  2. 2)
    This work pioneers the simultaneous optimization of two discrete variables, gear shift and clutch engagement/disengagement, using reinforcement learning for PHEV energy management. In contrast to previous studies [22] that rely on rule-based methods, this research emphasizes the systematic optimization [24] of gear shift and clutch engagement/disengagement for further fuel consumption reduction. The proposed approach is expected to yield superior energy management performance.
II Continuous-Discrete Reinforcement Learning
II-A Reinforcement Learning
@@
Reinforcement learning is a learning approach for optimal control by interacting with the environment. The objective is to develop an agent that can learn to make optimal decisions by maximizing a cumulative reward signal. Typically, this problem is formulated as a Markov decision process (MDP), which consists of a state space $\mathcal{S}$, an action space $\mathcal{A}$, an initial state distribution $p_{0}$, and a transition probability $p(s^{\prime},r|s,a)$ that specifies the probability of transitioning to state $s^{\prime}$ and receiving reward $r$ when taking action $a$ in state $s$, where $r(s,a)\in\mathbb{R}$. The agent’s goal is to maximize the expected cumulative reward $\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$, where $\gamma\in[0,1]$ is the discount factor.
@@
II-B Mixed-Integer Action Space
In this study, we explore the mixed-integer action space from the perspective of a parallel architecture, where discrete and continuous actions are independent of each other. For instance, in a PHEV energy management problem, the gear shift command could be the discrete action, while the engine torque could be the continuous action. These actions can be executed simultaneously and separately. Such mixed-integer action spaces can be defined as follows:
$\mathcal{A}=\{a=(a^{d},a^{c})\mid a^{d}\in\mathcal{A}_{d},a^{c}\in\mathcal{A}_{c}\}$ (1)
where $\mathcal{A}_{d}\subseteq\mathbb{N}^{m}$ represents a finite set of discrete action spaces, $\mathcal{A}_{c}\subseteq\mathbb{R}^{n}$ denotes a continuous action space, and $a^{d}$ and $a^{c}$ are the discrete and continuous actions, respectively. The dimensionality of the action space is determined by $m+n$. To handle the action space for multi-dimensional discrete variables, a composite action vector can be employed. This vector has a length of $d_{1}\times d_{2}\times\cdots\times d_{m}$, where $d_{i}$ represents the number of discrete actions for the $i$-th dimension. Each element of the vector corresponds to a distinct combination of discrete actions.
II-C TD3AQ Algorithm
@@
To address the above mixed-integer action space, we propose a novel model-free algorithm, TD3AQ, which is based on the actor-critic framework. The forward propagation process of TD3AQ is shown in Fig. 1. The actor network $\mu_{\theta}$ is used to generate the continuous actions $a^{c}$, then $a^{c}$ and state $s$ are fed into the Q-network to obtain the Q-values $Q_{1},Q_{2},\ldots,Q_{k}$ for each discrete action. Finally, the discrete action with the highest Q-value is selected as $a^{d}$. The details of the TD3AQ algorithm are elaborated below.
II-C1 Deep Deterministic Actor-Q
@@
Deep deterministic actor-Q (DDAQ) [25] adopts the same architecture as DDPG, but instead of approximating a single Q-value, it approximates $k$ Q-values for the $k$ discrete actions. Therefore, the Bellman equation can be rewritten as
$Q_{\omega}(s,a)=\mathbb{E}_{r,s^{\prime}}\left[r+\gamma\max_{a^{d}\in\mathcal{A}_{d}}Q_{\omega}\left(s^{\prime},a^{d},\mu_{\theta}(s^{\prime})\right)\right]$ (6)
@@
When updating the actor network, since the discrete action with the highest Q-value is selected as the output, the objective function is defined as
@@
where the maximum Q-value is used as the objective function. In previous work (P-DQN) [15], the objective function was defined as the sum of all Q-values, which posed a problem when back propagating gradients: updating the weights in the Q-network during back propagation affected all weights in the network, including those unrelated to the action with the highest Q-value, potentially causing the algorithm to update in an incorrect direction. DDAQ addresses this issue by using the maximum Q-value as the objective function, ensuring that only the weights corresponding to the action $a^{d}$ with the highest Q-value are updated.
II-C2 TD3AQ
DDPG has shown potential in achieving good performance, but its robustness is limited due to its sensitivity to hyperparameters. This is mainly because DDPG tends to overestimate Q-values, which results in poor policy updates. To address this issue, we combine DDAQ and TD3 [26] to form TD3AQ. Three key modifications to improve the performance are introduced.
Target policy smoothing
a noise sampled from a normal distribution is added to the action generated by the target actor network. This technique acts as a regularizer that prevents the actor network from converging to a suboptimal solution. Specifically, the action was smoothed by adding a clipped noise, defined as follows:
@@ II-C2 TD3AQ
Clipped double Q-learning
to mitigate the overestimation of Q-values, a twin Q-network is introduced to DDAQ. When updating the Q-network, the minimum of the two selected Q-values is used as the target value:
@@
Delayed updates
the actor network and target networks are updated less frequently than the Q-networks. This helps to stabilize the training process since slowing down the actor allows the critic to learn more accurate Q-values before letting it guide the actor. The networks are updated every two iterations as recommended in [26].
@@
Input: Gradient stepsizes $\{\alpha,\beta,\rho\}\geq 0$, exploration noise $\eta=\mathcal{N}(0,\delta^{2})$, exploration parameter $\varepsilon$, minibatch size $B$, empty replay buffer $\mathcal{B}$, target smooth noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$, target smooth noise bounds $\epsilon_{l}$, $\epsilon_{h}$, discount factor $\gamma$, policy update frequency $K$.
1 Initialize Q-network $Q_{\omega_{1}}$, $Q_{\omega_{2}}$, and actor network $\mu_{\theta}$ with random weights $\omega_{1}$, $\omega_{2}$, $\theta$;
2 Initialize target networks $\omega_{1}^{\prime}\leftarrow\omega_{1}$, $\omega_{2}^{\prime}\leftarrow\omega_{2}$, $\theta^{\prime}\leftarrow\theta$;
3 Initialize replay buffer $\mathcal{B}$;
4 for $t=1,2,\ldots,T$ do

### v3

# Mixed-Integer Optimal Control via Reinforcement Learning: A Case Study on Hybrid Electric Vehicle Energy Management

@@ Abstract
Many optimal control problems require the simultaneous output of discrete and continuous control variables. These problems are usually formulated as mixed-integer optimal control (MIOC) problems, which are challenging to solve due to the complexity of the solution space. Numerical methods such as branch-and-bound are computationally expensive and undesirable for real-time control. This paper proposes a novel hybrid-action reinforcement learning (HARL) algorithm, twin delayed deep deterministic actor-Q (TD3AQ), for MIOC problems. TD3AQ combines the advantages of both actor-critic and Q-learning methods, and can handle the discrete and continuous action spaces simultaneously. The proposed algorithm is evaluated on a plug-in hybrid electric vehicle (PHEV) energy management problem, where real-time control of the discrete variables, clutch engagement/disengagement and gear shift, and continuous variable, engine torque, is essential to maximize fuel economy while satisfying driving constraints. Simulation outcomes demonstrate that TD3AQ achieves control results close to optimality when compared with dynamic programming (DP), with just 4.69% difference. Furthermore, it surpasses the performance of baseline reinforcement learning algorithms.
keywords:
@@
mixed-integer optimal control , reinforcement learning , hybrid action space , hybrid electric vehicle energy management
journal: Engineering Applications of Artificial Intelligence
[…]
1 Introduction
Mixed-integer optimal control (MIOC) problems commonly arise in various engineering applications, such as assignment problems [1], control of piecewise-affine (PWA) systems [2] and hybrid systems [3]. These problems are particularly challenging as they require optimization of a system that involves both continuous and discrete control variables. Traditional numerical optimization methods, such as branch-and-bound and heuristic algorithms, are computationally expensive or rely on system models [4, 5, 6], making them unpreferred for general real-time control applications. Recent advances in reinforcement learning (RL) have demonstrated promising results in handling complex control problems [7, 8, 9]. However, most existing RL algorithms are tailored for either continuous [10] or discrete [7] action spaces, and are not directly applicable to MIOC problems. Currently, research on MIOC through RL is still in its nascent stages. This paper proposes an innovative hybrid-action reinforcement learning (HARL) algorithm for MIOC problems, and applies it to a plug-in hybrid electric vehicle (PHEV) energy management problem to validate its effectiveness.
1.1 Literature Review
Various methods have been proposed to solve MIOC problems. However, the unique characteristics of these methods still present certain limitations.
One approach is to formulate MIOC problems as mixed-integer programming (MIP) forms and solve them using the branch-and-bound method [11]. This method typically involves dividing the problem into smaller sub-problems or branches, and then solving each branch separately. A bounding function is used to determine if each branch can be pruned or eliminated, until the optimal solution is found. While this method can achieve globally optimal solutions, it is computationally expensive and time-consuming.
Another approach is to use dynamic programming (DP), which breaks the problem into smaller sub-problems and solves them recursively [12]. However, DP cannot handle continuous variables naturally and requires discretization of the continuous variables into a finite set of discrete values [13]. A fine-grained discretization can lead to a more accurate solution, but it also increases the computational cost, making it unsuitable for large-scale problems. Additionally, DP is not suitable for real-world optimal control since it requires knowledge of the entire state space in advance.
Heuristic algorithms, while not providing optimality guarantees, can offer feasible and fast solutions to MIOC problems [5, 6]. These methods use problem-specific knowledge to guide the search for a solution, rather than exhaustively examining every possible solution. They are simple to design and can be applied to a range of combinatorial optimization problems. However, it requires specific implementation and modification to match the problem structure being solved [6].
@@ 1.1 Literature Review
In recent years, a substantial amount of RL research has focused on a special type of MIOC problems called parameterized action Markov decision process (PAMDP) [17]. In PAMDP, the core idea involves structuring the problem hierarchically, with each discrete action linked to its own continuous parameters. One approach is to alternate between optimizing the discrete action selection with Q-learning and the parameter selection with a policy search method [17]. Another approach called parameterized deep deterministic policy gradient (PA-DDPG) is to directly output the continuous parameters and weights for discrete actions from DDPG [10] and select the discrete action with the highest weight [18]. Parameterized deep Q-networks (P-DQN) combine the actor-critic architecture and DQN, where the actor network first generates parameters for every discrete action, followed by the Q-network’s selection of the action that offers the maximum Q-value [19]. However, the formulation of P-DQN is flawed because the discrete action values depend on all continuous action parameters, not just those associated with each action. Therefore, Bester et al. [20] proposed a multi-pass DQN (MP-DQN) that used multiple forward passes from the actor network to the Q-network. Each pass involves only one set of action parameters for a discrete action.
In addition to the learning-based framework, Massaroli et al. [21] utilized neural ordinary differential equations as state-value networks to provide optimal action parameters for all discrete actions, then obtained the optimal discrete action using a value-based method. Li et al. [22] constructed the latent representation space of discrete actions and continuous parameters using a conditional variational auto-encoder. The RL agent’s output of the action and its parameters are initially fed into the representation model and subsequently exported to the environment. Overall, these methods may not be best for general MIOC problems as they rely on the parameterized action space.
Alternatively to PAMDP, some methods handle the mixed-integer action space in a unified manner. Jiang et al. [23] proposed a distributed dueling framework based on DQN, which added an extra fully-connected layer to generate continuous actions prior to the advantage and value layers. The discrete action was then selected to maximize the Q-value. Fan et al. [24] utilized a hybrid actor-critic architecture, two parallel actor networks were used to generate continuous and discrete actions, respectively. They shared the first few layers to encode the state information. Additionally, Neunert et al. [25] considered a hybrid policy that multiply the continuous and discrete actions distribution jointly, then optimized the policy with the maximum a posteriori policy optimization.
At present, there is limited research on RL for MIOC problems in engineering applications. In unmanned aerial vehicles control, Samir et al. [26] used greedy-based and DDPG-based methods to update discrete scheduling decision and continuous traveling distance and direction, respectively. Kamruzzaman et al. [27] adapted the architecture proposed by Fan et al. [24] to a multi-agent framework for installation planning and controlling of shunts to enhance power system resilience. Zhang et al. [28] and Ran et al. [29] adopted a structure similar to P-DQN [19] for MIOC in mobile edge computing systems and data center scheduling and cooling systems, respectively. Nonetheless, they utilized the maximum Q-value as the target Q-value for the policy, as opposed to the sum of all Q-values in P-DQN. Their approach is referred to as deep deterministic actor-Q (DDAQ), and it serves as the foundation for our proposed TD3AQ algorithm in this paper.
For energy management in HEVs, the primary objective of the optimal control strategy is to determine the torque distribution that effectively balances fuel efficiency and driving performance [30, 31]. Some research [32, 33] utilized a similar framework to Hausknecht et al. [18] to obtain the discrete driving mode using the $\arg\max$ function. Zhang et al. [34] treated the engine state (ON/OFF) as a binary classification problem and determined the discrete action output by a threshold. Tang et al. [35] integrated DQN and DDPG techniques to separately derive values for throttle settings and gear ratios. Similarly, Wang et al. [36] employed an approach akin to that of P-DQN [19], focusing on the regulation of engine torque and clutch states. Gong et al. [37] further extended the P-DQN algorithm by integrating a twin-delayed deep deterministic policy gradient (TD3) [38] algorithm to better handle the hybrid action space involved. These studies either simplify the problem to a single action space or consider the problem as a PAMDP. None of them solve the HEV energy management problem from a unified perspective.
@@
Inspired by the aforementioned research, this paper proposes a novel hybrid-action reinforcement learning algorithm called twin delayed deep deterministic actor-Q (TD3AQ) for MIOC problems. The algorithm is applied to address an energy management problem for a plug-in hybrid electric vehicle, aiming to demonstrate its effectiveness. The main contributions of this study are summarized as follows:
  • A novel HARL algorithm, TD3AQ, is proposed to address the mixed-integer action space. TD3AQ combines the advantages of both TD3 and Q-learning methods, providing a unified solution for MIOC problems while stabilizing the training process..
  • This study pioneers a comparative analysis of several prevalent RL algorithms within the domain of vehicular control applications, specifically focusing on PHEV energy management. Algorithms examined include DDAQ [29], PATD3 [18], and Rainbow [16] (discretized version). The evaluation results underscore TD3AQ’s capability to achieve near-optimal solutions compared to dynamic programming, surpassing the performance of prior HARL baselines.
The structure of the remainder of this paper is outlined below. Section 2 introduces the TD3AQ algorithm proposed in this work. Section 3 provides an overview of the PHEV powertrain model and elucidates the energy management problem. Section 4 presents the experimental results and offers insightful discussions. Finally, in Section 5, the paper is concluded, and avenues for future research are discussed.
2 Twin Delayed Deep Deterministic Actor-Q
2.1 Reinforcement Learning
@@
Reinforcement learning is a learning approach for optimal control by interacting with the environment. The objective is to develop an agent that can learn to make optimal decisions by maximizing a cumulative reward signal. Typically, this problem is formulated as a Markov decision process (MDP), which consists of a state space $\mathcal{S}$, an action space $\mathcal{A}$, an initial state distribution $p_{0}$, and a transition probability $p(s^{\prime},r|s,a)$ that specifies the probability of transitioning to state $s^{\prime}$ and receiving reward $r$ when taking action $a$ in state $s$, where $r(s,a)\in\mathbb{R}$. The agent’s goal is to maximize the expected cumulative reward $\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$, where $\gamma$ is a discount factor.
@@
2.2 Mixed-Integer Action Space
In this study, we explore the mixed-integer action space from the perspective of a parallel architecture, where discrete and continuous actions are independent of each other. For instance, in a PHEV energy management problem, the gear shift command could be the discrete action, while the engine torque could be the continuous action. These actions can be executed simultaneously and separately. Such mixed-integer action spaces can be described by
$\mathcal{A}=\{a=(a^{d},a^{c})\mid a^{d}\in\mathcal{A}_{d},a^{c}\in\mathcal{A}_{c}\}.$ (1)
In this formulation, $\mathcal{A}_{d}$, a subset of $\mathbb{N}^{k}$, constitutes the discrete action spaces, while $\mathcal{A}_{c}$, a subset of $\mathbb{R}^{n}$, comprises the continuous action spaces. Here, $a^{d}$ and $a^{c}$ respectively signify the actions within these discrete and continuous domains. The dimensionality of the action space is determined by $k+n$. To handle the action space for multi-dimensional discrete variables, a composite action vector can be employed. This vector has a length of $d_{1}\times d_{2}\times\cdots\times d_{k}$, where $d_{i}$ represents the number of discrete actions for the $i$-th dimension. Each element of the vector corresponds to a distinct combination of discrete actions.
2.3 TD3AQ Algorithm
@@
To address the above mixed-integer action space, we propose a novel model-free algorithm, TD3AQ, which is based on the actor-critic framework. The forward propagation process of TD3AQ is shown in Fig. 1. The actor network $\mu_{\theta}$ is used to generate the continuous actions $a^{c}$, then $a^{c}$ and state $s$ are fed into the Q-network to obtain the Q-values $Q_{1},Q_{2},\ldots,Q_{k}$ for each discrete action. Finally, the discrete action with the highest Q-value is selected as $a^{d}$. The details of the TD3AQ algorithm are elaborated below. We first introduce the DDAQ algorithm [28, 29], which is a variant of DDPG that can handle the mixed-integer action space, and then we present the TD3AQ algorithm, which combines the DDAQ algorithm with the TD3 algorithm [38].
2.3.1 Deep Deterministic Actor-Q
@@
DDAQ adopts the same architecture as DDPG, but instead of approximating a single Q-value, it approximates $k$ Q-values for the $k$ discrete actions. Therefore, the Bellman equation can be rewritten as
$Q_{\omega}(s,a)$ $=\mathbb{E}_{r,s^{\prime}}\left[r+\gamma\max_{a^{d}\in\mathcal{A}_{d}}Q_{\omega}\left(s^{\prime},a^{d},a^{c}\right)\right]$ (6)
$=\mathbb{E}_{r,s^{\prime}}\left[r+\gamma\max_{a^{d}\in\mathcal{A}_{d}}\left(Q_{\omega}\left(s^{\prime},a^{d},\mu_{\theta}(s^{\prime})\right)\right)\right]$ (7)
@@
Figure 2: The back propagation process of DDAQ. Assuming that the $Q_{i}$ has the highest Q-value, DDAQ only updates the parameters pertaining to action $a^{d}=i$ (the solid green lines in the figure).
During the update process of the actor network, the objective function is defined based on the selection of the discrete action that exhibits the highest Q-value:
@@
where the maximum Q-value is used as the objective function. In previous work (P-DQN) [19], the objective function was defined as the sum of all Q-values, which posed a problem when back propagating gradients: updating the weights in the Q-network during back propagation affected all weights in the network, including those unrelated to the action with the highest Q-value, potentially causing the algorithm to update in an incorrect direction. DDAQ addresses this issue by using the maximum Q-value as the objective function, ensuring that only the weights corresponding to the action $a^{d}$ with the highest Q-value are updated. The back propagation process of DDAQ is illustrated in Fig. 2.
2.3.2 TD3AQ
DDPG has shown potential in achieving good performance, but its robustness is limited due to its sensitivity to hyperparameters. This is mainly because DDPG tends to overestimate Q-values, which results in poor policy updates. To address this issue, we combine DDAQ and TD3 to form TD3AQ. Three key modifications to improve the performance are introduced.
1) Target policy smoothing: A noise sampled from a normal distribution is added to the action generated by the target actor network. This technique acts as a regularizer that prevents the actor network from converging to a suboptimal solution. Specifically, the action was smoothed by adding a clipped noise, defined as follows:
@@ 2.3.2 TD3AQ
2) Clipped double Q-learning: To mitigate the overestimation of Q-values, a twin Q-network is introduced to DDAQ. When updating the Q-network, the target value is set as the minimum of the two chosen Q-values:
@@
3) Delayed updates: The actor network and target networks are updated less frequently than the Q-networks. This helps to stabilize the training process since slowing down the actor allows the critic to learn more accurate Q-values before letting it guide the actor. The networks are updated every two iterations as recommended in [38].
@@
Input: Exploration noise $\eta=\mathcal{N}(0,\delta^{2})$, exploration parameter $\varepsilon$, minibatch size $B$, empty replay buffer $\mathcal{B}$, target smooth noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$, target smooth noise bounds $\epsilon_{l}$, $\epsilon_{h}$, discount factor $\gamma$, policy update frequency $K$, learning rates $\{\alpha,\beta,\rho\}\geq 0$.
1
2Set up Q-network $Q_{\omega_{1}}$, $Q_{\omega_{2}}$, and actor network $\mu_{\theta}$ with initial random weights $\omega_{1}$, $\omega_{2}$, $\theta$;
3
4Set up target networks $\omega_{1}^{\prime}\leftarrow\omega_{1}$, $\omega_{2}^{\prime}\leftarrow\omega_{2}$, $\theta^{\prime}\leftarrow\theta$;
5
6Set up replay buffer $\mathcal{B}$;
7
8for $t=1,2,\ldots,T$ do